OpenStudy (anonymous):

what is the equation of the line whose slope is 0 and that contains the point (-4, 2).

OpenStudy (lgbasallote):

0 slope means horizontal line

OpenStudy (mimi_x3):

\(y=mx+b\) \(m=0\) => \(x=-4,y=2\) Sub it in to find \(b\) After finding the variable \(b\) you will find the equation.

OpenStudy (anonymous):

okay you said a horizontal line?

OpenStudy (lgbasallote):

horizontal line yes

OpenStudy (anonymous):

so its either y+4=0 or y-2+0? at igbasallote?

OpenStudy (mimi_x3):

you don't need to worry about horzontal line i suppose.

OpenStudy (mimi_x3):

\(2 = 0(-4)+b\) => \( 2 = 0+b\) => \(b=2\) \(y= 0(x) +2\) => \(y=2\)

OpenStudy (anonymous):

so y-2=0?

OpenStudy (mimi_x3):

well.. \(y-2=0\) => \(y=2\) same equation.

OpenStudy (anonymous):

okay thanks!

OpenStudy (mimi_x3):

np